# On Phase Unwrapping via Digital Wavefront Sensors

### v1

@@ 4 Numerical experiments
In this section, we numerically evaluate our algorithms introduced in Section 3 on a real-world phase unwrapping problem appearing in free-space optical communication (FSOC), and compare the results to those obtained with state-of-the-art algorithms.
In FSOC, laser beams are used for the wireless transmission of data in e.g., optical laser communications between satellites and ground-based stations. Thanks to high bandwidths, laser communication provides large and fast data transfer rates. However, when the laser signal travels through the Earths’ atmosphere, it gets disrupted by turbulence. Recent research in FSOC, see e.g. [31, 8, 47, 38, 50], focuses on using Adaptive Optics (AO) to overcome this problem. AO systems have been employed for ground-based telescopes to compensate atmospheric distortions and to obtain diffraction-limited astronomical observations. In the design phase of AO systems for FSOC, simulations of laser light propagation including diffraction and interference are required. To this end, physical optics propagation is commonly employed [3], which uses Fourier transforms to compute phase angles. As these phase angles contain $2\pi$ ambiguities, fast and accurate phase unwrapping algorithms are of great interest for solving this problem.
@@ 4.1 Implementation and algorithms
The digital WFS phase unwrapping algorithms described in Section 3 were all implemented in MATLAB. As mentioned there, for the digital SH-WFS based unwrapping the CuReD method is used in Step 4 of Algorithm 2. This algorithm requires neither parameter tuning nor an initial guess [57, 43, 42]. Similarly, for Step 2 in Algorithm 4 the PCuReD algorithm is used. Finally, in Step 2 of Algorithm 3 the NOPE algorithm is used, which does require an initial guess, for which either zero, or an approximation computed via PCuReD can be used [21]. This is denoted below by “zero starting” and “linear starting”, respectively. Furthermore, NOPE also has an optional smoothness parameter $s$, which, unless noted otherwise, is set to the canonical choice $s=11/6$ as suggested in [21]; see also [25]. In order to validate the performance of our proposed algorithms, we compare them to a number of commonly used local and global phase unwrapping approaches. Among the local methods we use both the built-in MATLAB unwrapping function, as well as the more advanced algorithm proposed in [18]. The latter is based on unwrapping most reliable pixels (MRP), and has a MATLAB implementation available in [11]. Moreover, we use two global methods based on solving the Poisson equation [55] and the transport of intensity (TIE) equation [56], respectively. The corresponding MATLAB codes are available in [58] and [59].
@@ 30 $\times$ 30 100 $\times$ 100 500 $\times$ 500
Figure 4.2 shows the unwrapped phase $\phi_{r}$ obtained via the digital SH-WFS phase unwrapping approach, i.e., Algorithm 2, for three different choices of the number of subapertures/subdomains $\Omega_{j,k}$, as well as the corresponding reconstruction error. Compared to the ground truth phase $\phi$ depicted in Figure 4.1, the best results are obtained with $100\times 100$ subapertures. In the case of $30\times 30$ subapertures, the structure of the true phase $\phi$ is also still clearly visible, but the resolution is limited due to the comparatively low number of subdomains. However, in the case of $500\times 500$ subapertures the reconstruction is almost entirely corrupted by noise. This is because the number of subdomains implicitly acts as a regularization parameter in this approach, and thus has to be chosen appropriately to strike a proper balance between resolution and stability. Additional regularization may also be used, either by adapting Step 4 of Algorithm 2, or for example via the averaged digital SH-WFS approach introduced in [19].
@@
Figure 4.3: Unwrapped phases $\phi_{r}$ obtained via the different Fourier-based WFS unwrapping methods of Section 3.4, i.e., Algorithm 3 with the different functions $\psi$ listed in Table 3.1, applied to the noisy wrapped phase $\phi_{w}^{\delta}$ depicted in Figure 4.1 (top), as well as the corresponding absolute difference to the non-wrapped phase $\phi$ (bottom).
Next, we consider the results obtained via the digital Fourier-based WFS unwrapping methods of Section 3.4, in particular Algorithm 3 and Algorithm 4. We start with the classical 4-sided PWFS, cf. Table 3.1. Figure 4.3 collects four different reconstructions: The left column was obtained using the linear PCuReD method, i.e. Algorithm 4, while the other columns were obtained via the NOPE algorithm, i.e., Algorithm 3. As explained above, the smoothness index $s$ is an optional parameter of NOPE, and the terms zero/linear starting correspond to different initial guesses for the iterative procedure underlying NOPE [21]. In particular, the reconstruction shown in the left column was used as an initial guess for the NOPE reconstructions depicted in the right two columns. As can be seen from the residual plots, the best reconstruction is obtained when NOPE is combined with linear starting and $s=11/6$, as depicted in the third column. Good results are also obtained with the linear Algorithm 4 and Algorithm 3 with $s=1/2$, depicted in the first and fourth columns, respectively. Compared with the results shown in the second column, the benefit of linear starting over zero starting becomes apparent. Unfortunately, while the NOPE method (Algorithm 3) produces some of the best results overall, it is also computationally expensive, requiring several minutes of runtime compared to a couple of seconds for all other algorithms considered in this paper. As a remedy, NOPE can be run in an abridged form using only a couple of iterations and linear starting. The obtained results are then very close to those shown in the third column, but with a significantly improved computation time.
@@
Figure 4.4: Unwrapped phases $\phi_{r}$ obtained via the different Fourier-based WFS unwrapping methods of Section 3.4, i.e., Algorithm 3 with the different functions $\psi$ listed in Table 3.1, applied to the noisy wrapped phase $\phi_{w}^{\delta}$ depicted in Figure 4.1 (top), as well as the corresponding absolute difference to the non-wrapped phase $\phi$ (bottom).
Next, we consider the effect of different shape functions $\psi$ on the performance of Algorithm 3. In particular, we consider $\psi$ corresponding to the 3-sided pyramid, roof, cone, and iQuad wavefront sensors as given in Table 3.1. The obtained unwrapped phases $\phi_{r}$ and the corresponding reconstruction errors are depicted in Figure 4.4. Note that here “Roof” refers to $\psi$ being the sum of the corresponding x-Roof and y-Roof functions. Among the different sensors, only the 3-sided pyramid WFS, and to some extent also the roof WFS, still provide acceptable reconstructions. For the cone sensor, only traces of the structure of the true phase are visible, hidden behind severe artefacts, which for the iQuad sensor dominate the reconstruction entirely. These artefacts may be explained with non-uniqueness issues that have already been observed in [9] for the iQuad sensor, and might be removed with further developments of the NOPE algorithm.
@@
Finally, we compare the results of our proposed digital WFS unwrapping approaches with those obtained by four different state-of-the art methods introduced above. The corresponding results are depicted in Figure 4.5. First, note that the built-in Matlab unwrapping function fails entirely, while the MRP method produces discontinuous, patch-like reconstructions which are also far off from the true phase. The best reconstruction is the one obtained via the PE method, which is similar to, but less detailed than that of the digital SH-WFS unwrapping. The TIE unwrapping captures many of the structural features of the ground truth phase, but some features almost appear to be inverted with respect to the ground truth. In summary, we find that our proposed digital SH-WFS and Fourier-based WFS methods perform at least as good as, and often much better than, these four standard approaches. In terms of reconstruction quality, the digital SH-WFS with a suitable number of subapertures, and the digital 4-sided PWFS using NOPE with linear starting produce the best results among our methods.
@@
For the second example, we consider a phase unwrapping problem from FSOC, in particular the real-world ALASCA (Advanced Laser Guide Star Adaptive Optics for Satellite Communication Assessments) project [1] currently under development, for which we now provide some brief background. Within the ALASCA project, an optical communication channel from the optical ground station of the European Space Agency (ESA) in Tenerife to communication satellites is developed. The concept is based on laser guide star AO and aims for 24/7 operations. In the design phase of the project, reliable simulations of the laser guide star AO system are crucial. Hence, the laser propagation in a turbulent atmosphere has to be modelled, which usually involves physical optics propagation methods and thus phase screens that contain $2\pi$ ambiguities. The simulations for ALASCA are carried out via the PyrAmid Simulator Software for Adaptive opTics Arcetri (PASSATA)[2], which is IDL and CUDA based object oriented software for Monte-Carlo end-to-end AO simulations. Figure 4.6 shows a typical wrapped phase screen occurring during daytime conditions, i.e., strong turbulence, simulated with PASSATA, for which no ground truth unwrapped phase is available.
DSH Pyramid MATLAB
@@
Figure 4.7: Unwrapped phases $\phi_{r}$ obtained via different phase unwrapping approaches applied to the noisy wrapped phase $\phi_{w}^{\delta}$ depicted in Figure 4.6. From left top to right bottom: Digital SH-WFS unwrapping, i.e., Algorithm 2, digital 4-sided pyramid WFS unwrapping, i.e., Algorithm 4, MATLAB built-in function, unwrapping based on most reliable pixels (MRP), the Poisson equation (PE) and on the transport of intensity equation (TIE).
In Figure 4.7, we compare the unwrapped phases $\phi_{r}$ produced by the four state-of-the-art methods considered above to the two digital WFS approaches which performed best on simulated data: The digital SH-WFS, and the digital 4-sided PWFS using NOPE and linear starting. Due to the smaller size of the wrapped phase data, the shown results for the digital SH-WFS were obtained using $30\times 30$ subapertures. As before, the built-in MATLAB function and the MRP method produce unacceptable reconstructions, with either line or patch artefacts corrupting the phase. The other methods perform broadly similar, recovering the same general structure of the phase but with a varying amount of resolution. While due to the absence of a ground truth we are unable to say which of the unwrapping methods performs best, we nevertheless see that our digital WFS-based approaches produce unwrapped phases similar to those of the PE and TIE methods, and at a similar computational speed.
@@ 5 Conclusion
In this paper, we introduced a new class of phase unwrapping algorithms based on the concept of digital WFSs. In their derivation, the key observation was that when considered as wavefront aberrations in an optical system, wrapped and unwrapped phases produce the same sensor measurements. Since reconstruction algorithms for WFSs are typically optimized to provide smooth wavefront reconstructions, they can applied to compute an unwrapped phase from digitally computed WFS measurements. This procedure was then elaborated in detail for two classes of digital WFSs, based on both the SH and Fourier-based WFSs. Numerical examples on a real-world adaptive optics unwrapping problem demonstrate that the obtained results are comparable to, and in some cases exceed, the performance of other state-of-the-art unwrapping algorithms.

### v2

@@ 4 Numerical experiments
In this section, we numerically evaluate our algorithms introduced in Section 3 on a real-world phase unwrapping problem appearing in FSOC, and compare the results to those obtained with state-of-the-art algorithms.
In FSOC, laser beams are used for the wireless transmission of data in e.g. optical laser communications between satellites and ground-based stations. Thanks to high bandwidths, laser communication provides large and fast data transfer rates. However, when the laser signal travels through the Earth’s atmosphere, it gets disrupted by turbulence. Recent research in FSOC, see e.g. [31, 8, 47, 38, 50], focuses on using AO to overcome this problem. In the design phase of AO systems for FSOC, simulations of laser light propagation including diffraction and interference are required. To this end, physical optics propagation is commonly employed [3], which uses Fourier transforms to compute phase angles. As these phase angles contain $2\pi$ ambiguities, fast and accurate phase unwrapping algorithms are of great interest for solving this problem.
@@ 4.1 Implementation and algorithms
The digital WFS phase unwrapping algorithms described in Section 3 were all implemented in MATLAB. As mentioned above, for the digital SH-WFS based unwrapping the CuReD method is used in Step 4 of Algorithm 2. This algorithm requires neither parameter tuning nor an initial guess [57, 43, 42]. Similarly, for Step 2 in Algorithm 4 the PCuReD algorithm is used. Finally, in Step 2 of Algorithm 3 the NOPE algorithm is used, which does require an initial guess, for which either zero, or an approximation computed via PCuReD has been used [21]. This is denoted below by “zero starting” and “linear starting”, respectively. Furthermore, the NOPE also has an optional smoothness parameter $s$, which, unless noted otherwise, is set to the canonical choice $s=11/6$ as suggested in [21]; see also [25]. In order to validate the performance of our proposed algorithms, we compare them to a number of commonly used local and global phase unwrapping approaches. Among the local methods we use both the built-in MATLAB unwrapping function, as well as the more advanced algorithm proposed in [18]. The latter is based on unwrapping most reliable pixels (MRP), and has a MATLAB implementation available in [11]. Moreover, we use two global methods based on solving the Poisson equation [55] and the transport of intensity (TIE) equation [56], respectively. The corresponding MATLAB codes are available in [58] and [59].
@@ 30 $\times$ 30 100 $\times$ 100 500 $\times$ 500
Figure 4.2 shows the unwrapped phase $\phi_{r}$ obtained via the digital SH-WFS phase unwrapping approach, i.e., Algorithm 2, for three different choices of the number of subapertures/subdomains $\Omega_{j,k}$, as well as the corresponding reconstruction error. Compared to the ground truth phase $\phi$ depicted in Figure 4.1, the best results are obtained with $100\times 100$ subapertures. In the case of $30\times 30$ subapertures, the structure of the true phase $\phi$ is also still clearly visible, but the resolution is limited due to the comparatively low number of subdomains. However, in the case of $500\times 500$ subapertures the reconstruction is almost entirely corrupted by noise. This is because the number of subdomains implicitly acts as a regularization parameter in this approach. Thus it has to be chosen appropriately to strike a proper balance between resolution and stability. Additional regularization may also be used, either by adapting Step 4 of Algorithm 2, or for example via the averaged digital SH-WFS approach introduced in [19].
@@
Figure 4.3: Unwrapped phases $\phi_{r}$ obtained via the different PWFS unwrapping methods of Section 3.4, i.e., Algorithm 34, applied to the noisy wrapped phase $\phi_{w}^{\delta}$ depicted in Figure 4.1 (top), as well as the corresponding absolute difference to the non-wrapped phase $\phi$ (bottom). The results show Algorithm 4 (left) and Algorithm 3 (three right columns) with zero or linear starting and smoothness index $s=11/6$ or $s=1/2$ respectively.
Next, we consider the results obtained via the digital Fourier-type WFS unwrapping methods of Section 3.4, in particular Algorithm 3 and Algorithm 4. We start with the classical 4-sided PWFS, cf. Table 3.1. Figure 4.3 collects four different reconstructions: The left column was obtained using the linear PCuReD method, i.e. Algorithm 4, while the other columns were obtained via the NOPE algorithm, i.e., Algorithm 3. As explained above, the smoothness index $s$ is an optional parameter of the NOPE, and the terms zero/linear starting correspond to different initial guesses for the iterative procedure underlying the NOPE [21]. In particular, the reconstruction shown in the left column was used as an initial guess for the NOPE reconstructions depicted in the right two columns. As can be seen from the residual plots, the best reconstruction is obtained when the NOPE is combined with linear starting and $s=11/6$, as depicted in the third column. Good results are also obtained with the linear Algorithm 4 and Algorithm 3 with $s=1/2$, depicted in the first and fourth columns, respectively. Compared with the results shown in the second column, the benefit of linear starting over zero starting becomes apparent. Unfortunately, while the NOPE method (Algorithm 3) produces some of the best results overall, it is also computationally expensive, requiring several minutes of runtime compared to a couple of seconds for all other algorithms considered in this paper. As a remedy, the NOPE can be run in an abridged form using only a couple of iterations and linear starting. The obtained results are then very close to those shown in the third column, but with a significantly improved computation time.
@@
Figure 4.4: Unwrapped phases $\phi_{r}$ obtained via the different Fourier-type WFS unwrapping methods of Section 3.4, i.e., Algorithm 3 with the different functions $\psi$ listed in Table 3.1, applied to the noisy wrapped phase $\phi_{w}^{\delta}$ depicted in Figure 4.1 (top), as well as the corresponding absolute difference to the non-wrapped phase $\phi$ (bottom).
Next, we consider the effect of different shape functions $\psi$ on the performance of Algorithm 3. In particular, we consider $\psi$ corresponding to the 3-sided pyramid, roof, cone, and iQuad wavefront sensors as given in Table 3.1. The obtained unwrapped phases $\phi_{r}$ and the corresponding reconstruction errors are depicted in Figure 4.4. Note that here “Roof” refers to the averaged sum of the corresponding x-Roof and y-Roof reconstructions. Among the different sensors, only the 3-sided PWFS, and to some extent also the roof WFS, still provide acceptable reconstructions. For the cone sensor, only traces of the structure of the true phase are visible, hidden behind severe artefacts, which for the iQuad sensor dominate the reconstruction entirely. These artefacts may be explained with non-uniqueness issues that have already been observed in [9] for the iQuad sensor, and might be removed with further developments of the NOPE algorithm.
@@
Finally, we compare the results of our proposed digital WFS unwrapping approaches with those obtained by four different state-of-the art methods introduced above. The corresponding results are depicted in Figure 4.5. First, note that the built-in Matlab unwrapping function fails entirely, while the MRP method produces discontinuous, patch-like reconstructions which are also far off from the true phase. The best reconstruction is the one obtained via the PE method, which is similar to, but less detailed than that of the digital SH-WFS unwrapping. The TIE unwrapping captures many of the structural features of the ground truth phase, but some features almost appear to be inverted with respect to the ground truth. In summary, we find that our proposed digital SH-WFS and Fourier-type WFS methods perform at least as good as, and often much better than, these four standard approaches. In terms of reconstruction quality, the digital SH-WFS with a suitable number of subapertures, and the digital 4-sided PWFS using the NOPE with linear starting produce the best results among our methods.
@@
For the second example, we consider a phase unwrapping problem from FSOC, in particular the Advanced Laser Guide Star Adaptive Optics for Satellite Communication Assessments (ALASCA) project [1] currently under development, for which we now provide some brief background. Within the ALASCA project, an optical communication channel from the optical ground station of the European Space Agency (ESA) in Tenerife to communication satellites is developed. The concept is based on laser guide star AO and aims for 24/7 operations. In the design phase of the project, reliable simulations of the laser guide star AO system are crucial. Hence, the laser propagation in a turbulent atmosphere has to be modelled, which usually involves physical optics propagation methods and thus phase screens that contain $2\pi$ ambiguities. The simulations for ALASCA are carried out via the PyrAmid Simulator Software for Adaptive opTics Arcetri (PASSATA)[2], which is IDL and CUDA based object oriented software for Monte-Carlo end-to-end AO simulations. Figure 4.6 shows a typical wrapped phase screen occurring during daytime conditions, i.e., strong turbulence, simulated with PASSATA, for which no ground truth unwrapped phase is available.
SH-WFS PWFS + NOPE MATLAB
@@
Figure 4.7: Unwrapped phases $\phi_{r}$ obtained via different phase unwrapping approaches applied to the noisy wrapped phase $\phi_{w}^{\delta}$ depicted in Figure 4.6. From left top to right bottom: Digital SH-WFS unwrapping, i.e., Algorithm 2, digital 4-sided PWFS unwrapping, i.e., Algorithm 3, MATLAB built-in function, unwrapping based on most reliable pixels (MRP), the Poisson equation (PE) and on the transport of intensity equation (TIE).
In Figure 4.7, we compare the unwrapped phases $\phi_{r}$ produced by the four state-of-the-art methods considered above to the two digital WFS approaches which performed best on simulated data: The digital SH-WFS, and the digital 4-sided PWFS using the NOPE and linear starting. Due to the smaller size of the wrapped phase data, the shown results for the digital SH-WFS were obtained using $30\times 30$ subapertures. As before, the built-in MATLAB function and the MRP method produce unacceptable reconstructions, with either line or patch artefacts corrupting the phase. The other methods perform broadly similar, recovering the same general structure of the phase but with a varying amount of resolution. While due to the absence of a ground truth we are unable to say which of the unwrapping methods performs best, we nevertheless see that our digital WFS-based approaches produce unwrapped phases similar to those of the PE and TIE methods, and at a similar computational speed.
@@ 5 Conclusion
In this paper, we introduced a new class of phase unwrapping algorithms based on the concept of digital WFSs. In their derivation, the key observation was that when considered as wavefront aberrations in an optical system, wrapped and unwrapped phases produce the same sensor measurements. Since reconstruction algorithms for WFSs are typically optimized to provide smooth wavefront reconstructions, they can be applied to compute an unwrapped phase from digitally computed WFS measurements. This procedure was then elaborated in detail for two classes of digital WFSs, based on both the SH and Fourier-type WFSs. Numerical examples on a real-world adaptive optics unwrapping problem demonstrate that the obtained results are comparable to, and in some cases exceed, the performance of other state-of-the-art unwrapping algorithms.